# LATTE: Learning Aligned Transactions and Textual Embeddings for Bank Clients

### v1

@@ 4.2 Runtime Analysis
Figure 3 showcases the performance on the Churn dataset. Generative models (LATTE [1, 2, 3], TALLRec) achieve the highest ROC-AUC scores, exceeding 0.868, but at the cost of extremely low inference speeds—processing only a few samples per second—and high memory consumption, with over 3 billion parameters. In contrast, lightweight LATTE-S variants attain slightly lower ROC-AUC values while operating over 160 samples per second and compact size of just 4.4 million parameters. These models match the efficiency of contrastive baselines, offering a favorable trade-off between accuracy, speed, and memory footprint. As shown in Appendix B (Figure 4), a similar trend occurs on the Age prediction task.
4.3 Ablation Study
Contrastive Fine-tuning with Textual Embedding Alignment.
@@ 4.3 Ablation Study
Descriptions 0.718 ± 0.004 0.552 ± 0.002 0.692 ± 0.006
CoLES + descriptions 0.853 ± 0.009 0.662 ± 0.005 0.895 ± 0.008
LATTE-S[1] 0.852 ± 0.005 0.657 ± 0.007 0.890 ± 0.008
LATTE [1] 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE-S[2] 0.858 ± 0.007 0.662 ± 0.006 0.894 ± 0.009
LATTE [2] 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE-S[3] 0.844 ± 0.005 0.663 ± 0.004 0.899 ± 0.001
LATTE [3] 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
@@
Table 2 summarizes an ablation study on contrastive fine-tuning and modality concatenation. Adding LLM-generated descriptions consistently boosts performance across all heads—for instance, LATTE [2] improves from 0.858 to 0.868 AUC on churn and from 0.894 to 0.901 on gender. Contrastive alignment proves essential, as all LATTE variants outperform CoLES + descriptions, which rely only on feature concatenation. LATTE [3] achieves the best results on age (0.678) and gender (0.902), highlighting the impact of regularization. These results show that language grounding via contrastive supervision outperforms simple fusion.
@@
Contrastive Head LLM Generator Churn (AUC) Age Group (Acc) Gender (AUC)
LATTE [1] Gemma 3 4B 0.849 ± 0.005 0.645 ± 0.006 0.888 ± 0.005
LATTE [1] Gemma 3 27B 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [1] Qwen 3 32B 0.859 ± 0.004 0.658 ± 0.005 0.896 ± 0.003
LATTE [2] Gemma 3 4B 0.851 ± 0.005 0.652 ± 0.006 0.889 ± 0.002
LATTE [2] Gemma 3 27B 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [2] Qwen 3 32B 0.862 ± 0.001 0.667 ± 0.004 0.896 ± 0.000
LATTE [3] Gemma 3 4B 0.849 ± 0.005 0.665 ± 0.006 0.890 ± 0.005
LATTE [3] Gemma 3 27B 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
LATTE [3] Qwen 3 32B 0.862 ± 0.004 0.675 ± 0.002 0.895 ± 0.003
@@
Table 3 evaluates how the choice of LLM for behavioral description generation affects downstream task performance. Across all contrastive heads, larger instruction-tuned models consistently yield stronger results. In particular, Qwen 3 32B outperforms smaller alternatives on churn prediction across all heads, achieving the best score of 0.862 AUC under LATTE [1]. Meanwhile, Gemma 3 27B provides competitive performance on age and gender classification, producing the top scores under both LATTE [2] (0.671 / 0.901) and LATTE [3] (0.678 / 0.902). In contrast, the smallest model, Gemma 3 4B, lags across all settings. These results emphasize that the generative capacity of the LLM is crucial for high-quality representation learning.
@@
LATTE [1] mE5-large-instruct 0.860 ± 0.005 0.658 ± 0.008 0.895 ± 0.003
LATTE [1] Qwen3-Embedding-0.6B 0.862 ± 0.004 0.660 ± 0.007 0.897 ± 0.003
LATTE [1] Qwen3-Embedding-8B 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [2] mE5-large-instruct 0.860 ± 0.007 0.664 ± 0.006 0.894 ± 0.004
LATTE [2] Qwen3-Embedding-0.6B 0.863 ± 0.005 0.667 ± 0.005 0.896 ± 0.003
LATTE [2] Qwen3-Embedding-8B 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [3] mE5-large-instruct 0.850 ± 0.004 0.671 ± 0.003 0.897 ± 0.002
LATTE [3] Qwen3-Embedding-0.6B 0.852 ± 0.006 0.673 ± 0.001 0.899 ± 0.002
LATTE [3] Qwen3-Embedding-8B 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
@@
Table 4 investigates the impact of the frozen text encoder used to obtain $z_{i}^{\text{text}}$ on downstream task performance, where the resulting embeddings serve both for alignment and as a standalone source of world knowledge. We compare mE5-large-instruct [44] with two variants of Qwen3 [45]: a lightweight embedding model (0.6B) and a large one (8B). Across all contrastive heads, Qwen3-Embedding-8B consistently delivers the strongest results, outperforming mE5 by up to 1.0–1.3% on age and gender tasks and up to 0.7% on churn. Notably, under LATTE [3], it pushes age classification to 0.678 and gender AUC to 0.902, both state-of-the-art. Appendix B provides additional ablation comparing different description embedding strategies.
Input Format for Prompting.
Contrastive Head Prompt Format Churn (AUC) Age Group (Acc) Gender (AUC)
LATTE [1] Raw serialization 0.854 ± 0.003 0.653 ± 0.007 0.887 ± 0.004
LATTE [1] Stats 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [2] Raw serialization 0.855 ± 0.005 0.656 ± 0.007 0.886 ± 0.005
LATTE [2] Stats 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [3] Raw serialization 0.842 ± 0.007 0.662 ± 0.005 0.889 ± 0.003
LATTE [3] Stats 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
Table 5: Ablation: Input format for prompting the LLM raw transaction serialization vs. summary statistics.
Table 5 compares two formats for prompting the LLM: raw serialized transaction sequences and structured behavioral summaries. Across all contrastive heads, using statistical summaries consistently outperforms raw serialization. For instance, in the LATTE [3] setup, statistical prompts improve churn ROC-AUC by 1.5% and gender AUC by 1.3%. Appendix B provides additional ablation of the textual description effect compared to raw statistical features.
5 Conclusion
We presented a novel method for contrastive representation learning from event sequences that leverages synthetically generated textual descriptions as a complementary modality. By aligning structured transaction data with natural language summaries produced by a frozen instruction-tuned LLM, the proposed approach introduces textual priors into the embedding space without requiring labeled data or LLM fine-tuning.
LATTE achieves state-of-the-art results across three key open-source banking tasks, with relative improvements of 6.1% in gender prediction, 3.0% in age group classification, and 2.0% in churn prediction compared to the baseline. The proposed LATTE-S is resource-efficient (4.4M parameters, up to 200 samples/sec speed), which is essential for industrial applications where behavioral logs are abundant but supervision is limited.
A particularly promising future direction is to explore richer forms of text-to-sequence alignment, where natural language summaries are more tightly coupled with the underlying event dynamics. This approach could yield embeddings that are not only effective for predictive tasks but also inherently interpretable, as they remain anchored in human-readable descriptions of user behavior.
Limitations
While our approach demonstrates strong empirical performance, it relies on the assumption that LLM-generated textual descriptions accurately reflect behavioral nuances present in structured sequences. Thus, the prompt design and generalization capacity of the frozen LLM limit the quality of supervision. Additionally, the text encoder is not updated during training, which may limit alignment fidelity in domains with distributional shifts. Finally, although our method does not require labels or fine-tuning, it introduces moderate training overhead compared to lightweight contrastive learning methods due to reliance on LLM-based generation at scale.
Appendix A Experimental Details
For all experiments, natural language descriptions were generated using the instruction-tuned Gemma-3-27B-Instruct [39] model. To encode the resulting behavioral descriptions into fixed-length vectors, we employed the Qwen3-Embedding-8B [40] model, a multilingual encoder optimized for semantic retrieval. The transaction encoder was instantiated as a GRU-based model trained under the CoLES self-supervised framework, serving as the base sequence encoder across all contrastive alignment heads. Both training and inference for the full pipeline, including LLM prompting, embedding computation, and contrastive alignment, were performed on a single NVIDIA Tesla A100 80GB GPU.
Appendix B Additional Experiments
Text Embedding vs. Stats Embedding
Contrastive Head Auxiliary Modality Churn (AUC) Age Group (Acc) Gender (AUC)
LATTE [1] Stats (MLP-1L) 0.846 ± 0.006 0.647 ± 0.006 0.886 ± 0.004
LATTE [1] Stats (MLP-2L) 0.847 ± 0.005 0.645 ± 0.007 0.887 ± 0.003
LATTE [1] LLM text 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [2] Stats (MLP-1L) 0.848 ± 0.005 0.651 ± 0.006 0.883 ± 0.005
LATTE [2] Stats (MLP-2L) 0.851 ± 0.003 0.653 ± 0.005 0.884 ± 0.004
LATTE [2] LLM text 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [3] Stats (MLP-1L) 0.842 ± 0.007 0.657 ± 0.004 0.886 ± 0.003
LATTE [3] Stats (MLP-2L) 0.843 ± 0.005 0.656 ± 0.003 0.887 ± 0.003
LATTE [3] LLM text 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
Table 6: Ablation: Effect of auxiliary modality used for alignment — handcrafted stats vs. LLM-generated text. Bold indicates best; underline second-best per group.
Table 6 compares two auxiliary supervision targets: raw statistical feature vectors (projected through MLPs) versus LLM-generated natural language descriptions. Text-based supervision leads to stronger downstream performance across all tasks for every contrastive head. The advantage is especially pronounced under LATTE [3], where LLM-based supervision improves age classification accuracy to 0.678 and gender AUC to 0.902 (the highest overall). Even the best-performing 2-layer MLP projections underperform by 1–2% in most settings. These results confirm that natural language descriptions offer more informative targets than handcrafted statistics for guiding sequence representation learning.
Impact of Text Embedding Extraction Strategy.
Contrastive Head Method Churn (AUC) Age Group (Acc) Gender (AUC)
LATTE [1] LLM encoder (ours) 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [1] MeanPool (generator) 0.859 ± 0.006 0.659 ± 0.003 0.893 ± 0.001
LATTE [2] LLM encoder (ours) 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [2] MeanPool (generator) 0.862 ± 0.005 0.666 ± 0.002 0.894 ± 0.003
LATTE [3] LLM encoder (ours) 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
LATTE [3] MeanPool (generator) 0.851 ± 0.007 0.671 ± 0.004 0.894 ± 0.002
We compare two strategies for obtaining text embeddings from behavioral descriptions. In our default setup, we use a frozen sentence encoder (Qwen3-Embedding-8B) to compute embeddings, separating the generation and encoding processes. As an alternative, we extract the embedding directly from the generator LLM (Gemma-3-27B) by mean-pooling hidden states. Following recent work suggesting that mean pooling over all token embeddings outperforms using the EOS token [46, 42], we average hidden states over the final $k=8$ transformer layers. For instance, employing a dedicated LLM encoder yields up to a 0.8% improvement in both gender and churn AUC over direct pooling from the generator.
@@ Appendix B Additional Experiments
Figure 4 presents results for the Age prediction task. We observe trends consistent with those in the Churn dataset (Figure 3): generative models attain the highest accuracy (up to 0.678) but exhibit limited inference throughput. In contrast, lightweight LATTE-S variants achieve a favorable trade-off, maintaining competitive performance ( 0.66 ROC-AUC) while delivering significantly higher inference speeds, exceeding 200 samples per second.

### v4

@@ 5.2 Runtime Analysis
Figure 3 showcases the performance on the Churn dataset. Generative models (LATTE, TALLRec) achieve the highest ROC-AUC scores, exceeding 0.868, but at the cost of extremely low inference speeds—processing only a few samples per second—and high memory consumption, with over 3 billion parameters. In contrast, lightweight LATTE-S variants attain slightly lower ROC-AUC values while operating over 160 samples per second and a compact size of just a few million parameters These models match the efficiency of contrastive baselines, offering a favorable trade-off between accuracy, speed, and memory footprint.
5.3 Contrastive Fine-tuning Alignment
@@
Descriptions 0.772 ± 0.011 0.432 ± 0.007 0.644 ± 0.009
$z^{\text{seq}}$ + $z^{\text{text}}$ 0.863 ± 0.007 0.650 ± 0.004 0.890 ± 0.003
LATTE-S[1] 0.847 ± 0.004 0.657 ± 0.003 0.891 ± 0.004
LATTE [1] 0.869 ± 0.004 0.665 ± 0.005 0.900 ± 0.005
LATTE-S[2] 0.846 ± 0.005 0.655 ± 0.004 0.888 ± 0.003
LATTE [2] 0.872 ± 0.004 0.663 ± 0.003 0.898 ± 0.006
@@
Table 2 presents an ablation study on the effect of contrastive fine-tuning and modality concatenation. Incorporating LLM-generated behavioral descriptions markedly improves downstream performance compared to CoLES, particularly for churn and age prediction tasks. The contrastive alignment step remains crucial: all LATTE variants outperform the non-aligned baseline. Among the evaluated heads, LATTE [2] achieves the highest AUC on churn (0.872), while LATTE [1] attains the best accuracy on age (0.665) and and gender (0.900). Nevertheless, unaligned concatenation remains a competitive baseline ($z^{\text{seq}}$ + $z^{\text{text}}$ ), indicating that the statistical–semantic descriptions alone already provide a strong inductive bias.
5.4 Evaluation of the quality of LLM summarization
Feature Churn Gender Age
mcc_0 Usage % 35.59 24.61 37.93
mcc_0 Acc % 100 100 100
mcc_1 Usage % 38.98 26.61 39.66
mcc_1 Acc % 100 100 100
trx_period Usage % 100 100 100
trx_period Acc % 98.31 99.02 99.14
trx_days_share Usage % 93.22 95.41 93.97
trx_days_share Acc % 98.18 92.38 99.07
Table 3: Four key LLM statistics usage (%) and accuracy (%) across tasks.
In this section, we study whether the textual descriptions generated by the LLM faithfully capture the underlying statistics that were used to construct the prompts. We asked an independent LLM critic (Llama 3.1 8B) to extract key statistics (e.g., dominant merchant categories, transaction period length, share of active days) from the subsample of generated descriptions. We then applied a rule-based matching procedure to compare these extracted factors against the ground-truth statistics.
Table 3 reports both the usage rate (how frequently a given statistic was mentioned in the LLM description) and the accuracy rate (the percentage of mentions that correctly reflect the underlying value) across three downstream tasks. We use 200 random samples per dataset. The results show that core statistics such as transaction_period and transaction_days_share are not only used very frequently (over 90% of cases), but also described with a high accuracy (above 92%). As expected, categorical statistics such as merchant-category (mcc_0, mcc_1) are consistently mentioned with 100% correctness when they appear.
Figure 4: UMAP visualizations of CoLES (left) and LATTE (right) embeddings colored by quantiles of different behavioral statistics. LATTE embeddings exhibit slightly better linear separation across some statistics such as (a) Total spending, (b) Total expense, (c) Average daily spending, and (d) Total spending in MCC (Eating places).
6 Analysis of Behavioral Embedding Structure
Metric CoLES LATTE
Total spending 0.627 0.693
Total expense 0.593 0.730
Average daily spending 0.519 0.689
Total spending in MCC (Eating places) 0.705 0.775
First transaction day 0.652 0.622
Std of daily spending 0.858 0.789
Table 4: Separability of user groups in the lowest 1% and highest 99% quantiles of behavioral statistics using logistic regression. LATTE embeddings show stronger separation across most statistics compared to CoLES.
In this section, we investigate how behavioral statistics are encoded within the embedding spaces of CoLES and LATTE (Figure 4). For each statistic, we highlight clients belonging to the lowest 1% (blue) and highest 99% (red) quantiles, and visualize the resulting structure using a UMAP projection with an overlaid logistic regression decision boundary. Compared to CoLES, LATTE embeddings exhibit more distinct geometric separation across representative behavioral dimensions-such as total spending, total expenses, average daily spending, and spending within the Eating places MCC category—indicating a stronger alignment between embedding geometry and behavioral variance.
To complement these qualitative observations, we conduct a quantitative assessment of separability for users with extreme behavioral profiles. As reported in Table 4, LATTE consistently achieves higher separability scores than CoLES, particularly for total expenses, average spending, and spending in top MCC categories, while performance remains comparable for total sum and weekday spending.
Figure 5: SHAP summary plot showing the most influential features contributing to model predictions.
Finally, the SHAP summary plot (Figure 5) provides further insight into feature importance within the predictive model. The most influential predictors include duration of activity, channel count, and average daily number of transactions, which show strong positive SHAP values for clients with higher feature magnitudes. Additionally, MCC-specific spending features and transaction-pattern embeddings make substantial contributions, implying that both aggregated behavioral indicators and categorical spending patterns play a crucial role in model predictions.
7 Conclusion
We presented a novel method (Fig. 2) for contrastive representation learning from event sequences that leverages synthetically generated textual descriptions as a complementary modality. By aligning structured transaction data with natural language summaries produced by a frozen instruction-tuned LLM, the proposed approach introduces textual priors into the embedding space without requiring labeled data or LLM fine-tuning. Our LATTE achieves state-of-the-art results across three key open-source banking tasks, with relative improvements of 6.1% in gender prediction, around 1.0% in age group classification, and 3.7% in churn prediction compared to the baseline. The proposed LATTE-S is resource-efficient (a few million parameters, up to 200 samples/sec speed), which is essential for industrial applications where behavioral logs are abundant but supervision is limited.
A particularly promising future direction is to explore richer forms of text-to-sequence alignment, where natural language summaries are coupled with the underlying event dynamics. This approach could yield to inherently interpretable embeddings.
Limitations
While our approach achieves strong empirical performance, it remains constrained by its reliance on a fixed set of pre-computed statistical features that condition the LLM-generated textual descriptions. This dependency limits adaptability when key behavioral patterns are not adequately captured by the chosen statistics. Furthermore, the method assumes that the generated descriptions faithfully reflect the underlying sequence dynamics, making performance sensitive to prompt design and the generalization capacity of the frozen LLM. Because the text encoder is not updated during training, alignment fidelity may further degrade under distributional shifts. Moreover, although the framework requires neither labels nor fine-tuning, it introduces moderate training overhead compared to lightweight contrastive objectives due to large-scale LLM-based generation.
Finally, in this paper, only financial transactions are used. The proposed sequence-to-text alignment framework can be extended to a wide range of domains that generate structured event logs-data types where LLMs often struggle due to sparsity, heterogeneity, and long temporal dependencies. Examples include healthcare Wang et al. (2024b), education Liu et al. (2023), e-commerce Dai et al. (2023); Liu et al. (2025). In these settings, LATTE could leverage LLM-generated descriptions to inject semantic priors into structural representations.
Acknowledgments
The work of A. Savchenko was supported by a grant, provided by the Ministry of Economic Development of the Russian Federation in accordance with the subsidy agreement (agreement identifier  [^00] C313925P4G0002) and the agreement with the Ivannikov Institute for System Programming of the Russian Academy of Sciences dated June 20, 2025 No. 139-15-2025-011.
Appendix A Experimental Details
Experimental Setup. For all experiments, natural language descriptions were generated using several instruction-tuned large language models of different types and scales, including Gemma-3-27B-Instruct, Gemma-2-27B-Instruct (Team et al., 2025), and Qwen3-Instruct-32B and Qwen3-Instruct-4B (Zhang et al., 2025). To encode the resulting behavioral descriptions into fixed-length vectors, we employed the Qwen3-Embedding-8B (Zhang et al., 2025) model, a multilingual encoder optimized for semantic retrieval. The transaction encoder was instantiated as a GRU-based model trained under the CoLES self-supervised framework, serving as the base sequence encoder across all contrastive alignment heads. Both training and inference for the full pipeline—including LLM prompting, embedding computation, and contrastive alignment—were performed on eight NVIDIA Tesla A100 80 GB GPUs.
Pipeline Details. A lightweight GRU-based transaction encoder was pretrained under the CoLES objective and later tuned via lightweight contrastive alignment heads, while keeping the text encoders frozen. Alignment was performed between sequence embeddings and combined textual–statistical embeddings derived from LLM-generated descriptions enriched with numerical features. For fair comparison across downstream models, feature selection was applied to all representations on the validation set, fixing the embedding dimensionality to 512 or 1024.
Appendix B Additional Experiments
Impact of Text Embedding Extraction Strategy.
Contrastive Head Method Churn (AUC) Age Group (Acc) Gender (AUC)
LATTE [1] LLM encoder 0.870 ± 0.005 0.665 ± 0.004 0.895 ± 0.002
LATTE [1] MeanPool 0.871 ± 0.003 0.663 ± 0.003 0.896 ± 0.001
We compare two strategies for obtaining text embeddings from behavioral descriptions. In our default setup, we use a frozen sentence encoder (Qwen3-Embedding-8B) to compute embeddings, separating the generation and encoding processes. As an alternative, we extract the embedding directly from the generator LLM (Gemma-3-27B) by mean-pooling hidden states. Following recent work suggesting that mean pooling over all token embeddings outperforms using the EOS token (BehnamGhader et al., 2024; Muennighoff et al., 2024), we average hidden states over the final $k=8$ transformer layers. Both versions provides similar embeddings quality.
@@ Appendix B Additional Experiments
Metric Gemma 3 4B Qwen 3 32B Gemma 3 27B
Churn (AUC) 0.872 ± 0.003 0.870 ± 0.005 0.869 ± 0.005
Age Group (Acc) 0.658 ± 0.007 0.665 ± 0.005 0.657 ± 0.004
Gender (AUC) 0.897 ± 0.005 0.895 ± 0.002 0.898 ± 0.003
@@
Table 6 examines how the choice of large language model for generating behavioral descriptions influences downstream performance. Among the compared generators, Qwen 3 32B achieves the highest average accuracy across tasks, leading on age group classification (0.665 Acc) and delivering competitive results on churn and gender prediction. Gemma 3 27B attains the best gender AUC (0.898) and remains comparable in other metrics, confirming that mid-scale models with strong instruction tuning can match much larger ones. In contrast, the compact Gemma 3 4B variant underperforms across all tasks. Overall, the results suggest that richer generative capacity and stronger instruction following in the LLM used for description synthesis translate into more informative and transferable sequence representations
@@
LATTE [1] mE5-large-instr 0.869 ± 0.005 0.662 ± 0.008 0.896 ± 0.002
LATTE [1] Qwen3-Emb-0.6B 0.862 ± 0.005 0.662 ± 0.004 0.899 ± 0.003
LATTE [1] Qwen3-Emb-8B 0.870 ± 0.005 0.665 ± 0.005 0.895 ± 0.002
@@
Table 7 analyzes how the choice of frozen text encoder for obtaining $z_{i}^{\text{text}}$ affects downstream task performance. We compare mE5-large-instruct (Wang et al., 2024a) with two Qwen3 embedding variants (Yang et al., 2025): a compact Qwen3-Emb-0.6B and a larger Qwen3-Emb-8B. The results show that all encoders perform comparably, with differences within a narrow margin of 0.5–1.0 pp across tasks. mE5-large-instruct yields the highest AUC on churn prediction (0.869), while Qwen3-Emb-8B slightly leads in age classification (0.665 Acc). The Qwen3-Emb-0.6B model achieves the best gender AUC (0.899), despite being the smallest.
@@
Figure 6 presents results for the Age prediction task. We observe trends consistent with those in the Churn dataset (Figure 3): generative models attain the highest accuracy (up to 0.665) but exhibit limited inference throughput. In contrast, lightweight LATTE-S variants achieve a favorable trade-off, maintaining competitive performance (0.57 ROC-AUC) while delivering significantly higher inference speeds, exceeding 200 samples per second.